OpenStudy (anonymous):

a chemist has 12 gallons of alcohol. he draws out 3 gallons and replaces it with water. he draws out 3 gallons of the mixture, again replacing it with water. after five such operations what is the amount of alcohol in gallons left in the mixture

OpenStudy (dumbcow):

ok first realize that total liquid is always 12 gallons first time it goes from 12 to 9 , then the ratio of alcohol to water is 9:12 or 3:4 2nd time of the 3 gallons taken out ,3/4 is alcohol or 2.25 gallons leaving 6.75 gal left 3rd time the ratio is 6.75:12 so 3*(6.75/12) or 1.6875 gallons of alcohol are taken out 4th time the ratio is 5.0625:12 ..... 1.265625 gal of alcohol taken out 5th time the ratio is 3.796875:12 .....0.94921875 gal of alcohol taken out Leaving 2.84765625 gallons of alcohol remaining

OpenStudy (dumbcow):

a more algebraic approach is to recognize the recursive relationship X_n+1 = X_n - 3(X_n/12) X_n+1 = (3/4)X_n this can be modeled by a geometric sequence X(n) = 12(3/4)^n plug in n=5 and you get X = 2.84765625

OpenStudy (dumbcow):

did that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how long have you been here?

OpenStudy (anonymous):

why tf does he need 12 gallons of alcohol?

OpenStudy (anonymous):

its a math question (idk)

OpenStudy (anonymous):

@dumbcow, are u still there?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

how long have you been on my question?

OpenStudy (anonymous):

i have one more question

OpenStudy (dumbcow):

oh like 5-10 min, i just happened to come back to this question after awhile

OpenStudy (anonymous):

in a certain population the ratio of the number if women to the number in men is 11-10. if the average (arithmetic mean) age of women is 34 and the average age of men s 32, the the average age of the population is A; 32 9/10 b;32 20/21 c; 33 d: 33 1/21 e; 33 1/10

OpenStudy (anonymous):

can you help with this one, my good sir?

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

what grade are u in

OpenStudy (anonymous):

im in 5th and in a really advanced class

OpenStudy (dumbcow):

wow you are really advanced for 5th grade im already done with school, im a private math tutor so i also help out on this site

OpenStudy (anonymous):

thats cool, except you have a really inappropriate picture for your image :)

OpenStudy (dumbcow):

avg = sum/N Lets say M is num of men, W is num of women ratio--> W/M = 11/10 --> W = 11/10M avgW = sumW/W = 34 --> sumW = 34*(11/10M) = 37.4M avgM = sumM /M = 32 --> sumM = 32M avg = (37.4M + 32M)/(M + 11/10M) = 33.05

OpenStudy (dumbcow):

oh i apologize about the pic, yeah it is inappropriate for young kids

OpenStudy (anonymous):

its okay (your image) just that you're a tutor and i wouldnt expect that image :)

OpenStudy (anonymous):

thanks for all of you help!

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

whats your name?